OpenStudy (anonymous):

Find the product: 1/3 x 6/25=??

OpenStudy (anonymous):

A. 1/25 B. 2/25 C. 7/25 D. 18/25

OpenStudy (anonymous):

\[\Large \frac{1}{3} \times \frac{6}{{25}} = \frac{6}{{75}} = \frac{?}{{25}}\]

OpenStudy (anonymous):

it's? B? maybe? idk?

OpenStudy (anonymous):

Well, what do you divide 75 by to get 25? Divide 6 by the same number (remember with fractions, whatever you do to the bottom, do to the top).